OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = -2x^2 - 5x + 7? x = 5/4 x = -2/7 x = 5 x = 7 <==my answer

OpenStudy (anonymous):

The normal to the tangent at the maximum value of the function would get you the axis of symmetry. Thus use second derivative to get the maximum value and then find its normal by using tangent. :)

jimthompson5910 (jim_thompson5910):

for the equation y = ax^2 + bx + c, the equation for the axis of symmetry is x = -b/(2a)

OpenStudy (anonymous):

I dont know why, i am getting x = -5/4 and i am pretty sure its right. :/

OpenStudy (anonymous):

well if i use jim_thompson5910's formula, I get 5/4. Maybe you forgot that two negatives cancel each other out?

OpenStudy (anonymous):

Yes sorry i used second derivative without cancelling the negative. Sorry. It is 5/4

OpenStudy (anonymous):

okay thanks!

jimthompson5910 (jim_thompson5910):

in this case, a = -2, b = -5 x = -b/(2a) x = -(-5)/(2*(-2)) x = 5/(-4) x = -5/4 So the axis of symmetry is x = -5/4

OpenStudy (anonymous):

okay actually, i just looked, and I miss typed my answer. It IS -5/4. I forgot the - part haha -oops-

jimthompson5910 (jim_thompson5910):

That's ok. I was wondering if there was a typo. Hopefully this is all making sense now?

OpenStudy (anonymous):

very much so!

OpenStudy (anonymous):

Well. This is awkward. I wasn't wrong after all i guess :))

jimthompson5910 (jim_thompson5910):

I'm glad it is